OpenStudy (anonymous):

A cube has a surface area of one hundred fifty square yards. What is the length of a side?

OpenStudy (ybarrap):

Surface area (\(SA\)) of a cube $$ SA=l\times w\times 6 $$ Where \(l\) is the length and \(w\) is the width. But this is a cube, so \(l=w\) So, $$ SA=6l^2=150\\ \implies l=\sqrt{\cfrac{150}{6}}=?\text{ yards} $$ Just calculate and \(l\) will be your answer.

OpenStudy (anonymous):

so what's the answer